OpenStudy (anonymous):

Help!! Im a little stuck.. I need to decompose this into partial fractions:

OpenStudy (anonymous):

\[(2x^2-11x+5)/(x-3)(x^2+x-4)\]

hero (hero):

What course are you currently taking?

OpenStudy (anonymous):

This is Pre-Calc Hero

hero (hero):

Okay, just wondering.

OpenStudy (anonymous):

Then I use A/(x-3)+B(x^2+x-4)

hero (hero):

What textbook are you using for the course?

OpenStudy (anonymous):

Precalculus: Graphs and Models by Addison Wesley

hero (hero):

Thanks

OpenStudy (anonymous):

Its pretty old. 1997

hero (hero):

That's okay. Sometimes, old is better.

OpenStudy (anonymous):

Okay. I mean its probably all the same techniques, for the most part.

hero (hero):

What edition?

hero (hero):

Well, nevermind. I'll get get the latest edition.

OpenStudy (anonymous):

Ok I wasnt sure. I dont know if it says it for sure.

hero (hero):

If you have the book, it will say what edition it is. You just have to look, but okay.

hero (hero):

I have about 9 min left before it finishes downloading

OpenStudy (anonymous):

I really dont see it at all :O This is the exact book though http://www.abebooks.com/book-search/isbn/0201694425/hard-cover/

OpenStudy (anonymous):

Ok thats fine. Thanks

hero (hero):

If you want to show someone a book, use amazon.com

OpenStudy (anonymous):

ahh alright.

hero (hero):

Yes, that version is very old. It's not even available for download. It's the first edition apparently

hero (hero):

4 minutes left..

OpenStudy (anonymous):

ok

hero (hero):

If you have the book, is there any reason why you are not able to understand this concept?

OpenStudy (anonymous):

I get the concepts down sort of, I just end up with the wrong numbers. I either do a step too early or Im not doing it completely right. I just want to make sure I understand it

hero (hero):

Okay, I will have a look here in about 15 secs

OpenStudy (anonymous):

Ok Ill go through with what Im doing and you can maybe see what Im doing wrong

OpenStudy (anonymous):

Let me know when its done

hero (hero):

It might take me a bit to figure this out.

OpenStudy (anonymous):

Ok thats fine!

hero (hero):

Okay, I see something similar to this where it talks about setting up a decomposition template.

OpenStudy (anonymous):

Its in a lesson called Partial Fractions in my book

OpenStudy (anonymous):

That sounds right though

hero (hero):

\[\frac{2x^2 - 11x+5}{(x-3)(x^2+x-4)} = \frac{A}{x-3} + \frac{Bx + C}{x^2 + x -4}\]

hero (hero):

Maybe that's what you're looking for.

OpenStudy (anonymous):

So then its \[A(x^2+x-4)+(x-3)(Bx+C)=2x^2-11x+5\]

hero (hero):

What I posted is in your book so it is correct. If what you posted is the next step, then by all means, go right ahead, but maybe someone else can take a look at that for you.

OpenStudy (anonymous):

Ok thats cool. I needed that for sure so thanks a lot Hero